AmiQWK II

QWKMail Offline Message System

for the Commodore Amiga

Version 2.3 April 1, 1994.

by James Dawson

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992,1993 James Dawson. All rights reserved. Permission is granted to distribute copies of this manual, The ‘AmiQWK’ executable, and all related files as long as none of the files in the distribution are altered or omitted. No profit or other material gain may be realized for the distribution of this program except to cover the costs of distribution. ‘AmiQWK’ may not be included with any commercial product without prior written permission from James Dawson. ‘AmiQWK’ may be distributed in freely distributable software libraries including the Fred Fish library. Users who register ‘AmiQWK’ will be given a registration key number. This registration number and the keyfile that ‘AmiQWK’ creates (‘AmiQWK.KEY’) are NOT to be distributed.

‘AmigaGuide’, ‘Amigaguide.info’, ‘amigaguide.library’ \begin{matrix}c \\ ◯ \end{matrix} Copyright 1992 Commodore-Amiga,

Inc. All Rights Reserved. Reproduced and distributed under license from Commodore. ‘AMIGAGUIDE’ SOFTWARE IS PROVIDED "AS-IS" AND SUBJECT TO CHANGE; NO WAR-RANTIES ARE MADE. ALL USE IS AT YOUR OWN RISK. NO LIABILITY OR RESPONSI-BILITY IS ASSUMED. ‘ReqTools’ and ‘reqtools.library’ are Copyright \begin{matrix}c \\ ◯ \end{matrix} Nico François. All product names used in this document are trademarks of their respective owners.

1 Introduction

1.1 About this manual

This manual assumes that you have at least a basic knowledge of the Amiga operating system. If you are not familiar with the Amiga I, recommend reading the ’Basic Operations’ chapter of your AmigaDOS manual.

This manual was written in ‘Texinfo’, the GNU documentation format, and processed through ‘MakeGuide’ to generate ascii, ‘AmigaGuide’, TEX, and PostScript versions.

1.2 What is QWKMail?

QWKMail is an offline message format. Using QWKMail allows you to call a BBS, collect all waiting messages, transfer them to your computer, and read them after you have logged off of the BBS. You can then read them with a QWK compatible message reader such as ‘AmiQWK’. New messages can be edited with your favorite text editor. Some of the advantages of using QWKMail include:

• Collecting all waiting messages into a QWKMail packet usually takes only a fraction of the time it takes to read messages online, resulting in lower phone bills.

• Since your replies are edited offline, you can take as much time as you need to enter and reply to messages. You also have the advantage of using tools that may not be available when using BBS editors, such as spelling and grammar checkers.

• You can easily save or print messages for future reference.

The QWK format was created in 1987 by Mark ’Sparky’ Herring. It is based on Clark Development Corp.’s ‘PC-Board’ Version 12.0 BBS message base format. The QWKMail format is now supported by most BBS programs either internally or by an external ’door’ program.

1.3 What is AmiQWK II?

‘AmiQWK II’ is a full featured QWKMail offline message reader for the Commodore Amiga computer. My goals for ‘AmiQWK’ are (in no particular order)

• To be the most stable and compatible QWKMail reader available for the Commodore Amiga computer.

• To be the most full featured QWKMail reader for the Commodore Amiga computer.

• To take advantage of features found Release 2 and higher of the Amiga operating system.

• To be easy to use.

• To be highly configurable.

2 Legal Stuff

2.1 About Shareware

‘AmiQWK II’ is released under the concept of Shareware. This means that you are granted a license to use ‘AmiQWK II’ for a period of 30 days after the initial use. After 30 days you must either register the program or discontinue its use. See Section 2.2 [Registration], page 3, for more details.

2.2 Registration

To register ‘AmiQWK’, fill out the registration form (filename ‘Registration.form’) and send it along with your payment to the address on the registration form. You will receive a 16 digit registration key based on your name. Enter this number into the ’Register’ window of the configuration editor and ‘AmiQWK II’ will generate a keyfile (‘AmiQWK.KEY’) that will remove the shareware registration notices.

Recommended method of payment is a check drwan from a U.S. bank or a money order in US dollars. (I can no longer accept personal checks drawn from Canadian banks due to a change in bank policy.) Cash is accepted (US Funds only), however you send cash at your own risk.

AmiQWK users in New Zealand and Australia can register through my New Zealand distributor. See the registration form (filename ‘registration.form’ for more details.

The registration key will work with all minor revisions of ‘AmiQWK II’. (i.e. all versions up to, but not including, 3.0) When a new major revision becomes available, registered ‘AmiQWK II’ users will be eligible for reduced cost updates.

Most registrations are processed within 2 working days. If a network address is given, I will send your registration key via email in addition to postal mail.

Group registrations and site licenses are available, Contact the author for details. See Appendix C [Technical Support], page 32.

2.3 Warranty

THIS PRODUCT IS PROVIDED "AS-IS" WITHOUT WARRANTY OF ANY KIND. ALL RISKS AND DAMAGES, INCIDENTAL OR OTHERWISE, ARISING FROM THE USE, MIS-USE, OR INABILITY TO USE THIS PROGRAM ARE ENTIRELY THE RESPONSIBILITY OF THE USER. THE AUTHOR DOES NOT MAKE ANY GUARANTEES OR REPRESENTA-TIONS REGARDING THE CORRECTNESS, RELIABILITY, ACCURACY, CURRENTNESS, ETC. OF THIS PROGRAM. THE AUTHOR WILL NOT ACCEPT RESPONSIBILITY FOR ANY DAMAGE OR LOSSES RESULTING FORM THE USE, MISUSE, OR INABILITY TO USE THIS PRODUCT.

2.4 Copyright

The ‘AmiQWK II’ software and documentation are copyright \begin{matrix}c \\ ◯ \end{matrix} 1992,1993 Jim Dawson. All rights reserved.

   ‘AmigaGuide’, ‘AmigaGuide.info’, ‘amigaguide.library’ \begin{matrix}c \\ ◯ \end{matrix} Copyright 1992 Commodore-

Amiga, Inc. All Rights Reserved. Reproduced and distributed under license from Commodore.

‘ReqTools’ and ‘reqtools.library’ are Copyright \begin{matrix}c \\ ◯ \end{matrix} Nico François.

All product names used in this document are trademarks of their respective owners.

2.5 Distribution

Regardless of the method of distribution, ‘AmiQWK II’ is NOT and never has been public domain.

‘AmiQWK II’ may be distributed freely, providing the following criteria are met:

• Distribution is limited to the original shareware distribution only. None of the files in the ‘AmiQWK II’ distribution archive may be modified or omitted.

• No material gain may be realized from the distribution of ‘AmiQWK II’, other than a small amount to cover distribution materials.

• ‘AmiQWK II’ may be included in freely distributable software libraries, including the Fred Fish collection.

• ‘AmiQWK II’ may not be bundled with any hardware or software product without prior written consent from Jim Dawson.

• Registration numbers for ‘AmiQWK’, and the key file that ‘AmiQWK II’ generates from them (‘AmiQWK.KEY’) may not be distributed.

3 System Requirements

3.1 Hardware Requirements

‘AmiQWK’ requires the following hardware:

• Amiga computer.

• 512KB RAM. (1MB or more recommended.)

• Floppy drive (Second floppy drive or hard drive highly recommended.)

• Printer. (if you want to print messages.)

3.2 Software Requirements

‘AmiQWK’ requires the following software:

• Workbench 2.04 (V37) or higher.

• A text editor that can save ascii text without any formatting info. (ED and MEMACS that come with Workbench are fine.)

• Any archivers you plan to use. (‘ZIP’, ‘LHA/LHArc’, ‘ZOO’, and ‘ARC’ are currently supported)

• ‘reqtools.library’. (included with ‘AmiQWK II’ distribution archive.)

• ‘amigaguide.library’. (included with ‘AmiQWK II’ distribution archive, only required for online help.)

4 Installation and Quick Setup

To install ‘AmiQWK’, just double-click on the ’Install AQ2’ icon. This will copy the ‘reqtools.library’ and ‘amigaguide.library’ to your ‘libs:’ directory. If newer versions of either library are already present, the library will not be copied.

Once the libraries have been installed, run ‘AmiQWK’ by double-clicking on it’s icon or by typing ‘AmiQWK’ from the ‘CLI’.

Once ‘AmiQWK’ is running, you will need to configure it to work on your system. Click on the gadget labeled ’Config ‘AmiQWK’…’ or type ’C’^{1} to bring up the configuration editor window.

The configuration editor window has 9 gadgets, the first 6 of which when clicked on will bring up a configuration sub-window. The remaining 3 gadgets are used to save the configuration to disk, use the configuration without saving, and to cancel changes you have made to the configuration. The following describes only the most basic configuration options needed to get ‘AmiQWK’ running, See Chapter 7 [Configuration], page 23, for a more detailed description of all of the configuration options.

‘Operation’

Change the ’Editor’ string gadget to whatever editor you use. About the only restrictions on the editor are that it must save text in plain ascii format (with no formatting codes) and that it does not detach itself from the CLI process that spawned it. Here are some examples:

CygnusEd c:ed -sticky (See Section 7.3 [CygnusEd Notes], page 25.)

          TurboText
                    c:ttx WAIT
          Ed (AmigaDos)
                    c:ed
          MEmacs (AmigaDos)
                    sys:tools/memacs
          GoldEd    c:ed sticky

See Section 7.2 [Operation window], page 24, for more information.

‘Replies’

1 Most actions in ‘AmiQWK’ can be accomplished by using either the mouse or the keyboard. Whenever a keyboard shortcut exists it will usually have the shortcut key underlined.

The default configuration for replies should not need to be changed for most applications. See Section 7.4 [Replies window], page 26, for more information.

‘Paths’ In the Paths window you tell ‘AmiQWK’ where it should look for and store certain files.

‘Packet Dir’

This is where ‘AmiQWK’ should look for mail packets. This is normally the same directory your term program uses for downloads.

‘Reply Dir’

This is where ‘AmiQWK’ stores replies. This is normally the same directory your term program uses for uploads.

‘Work Dir’ This is where ‘AmiQWK’ creates its work directory. If you have enough memory, you might want to change this to ‘RAM:’ for best performance. If you leave it blank ‘AmiQWK’ will create the work directory in the ‘AmiQWK’ program directory.

See Section 7.5 [Paths window], page 27, for more information.

‘Archivers’

You will probably not need to change anything in the Archivers window. See Section 7.6 [Archivers window], page 27, for more information.

‘Graphics’

You will probably not need to make changes in the Graphics window to get ‘AmiQWK’ to work. See Section 7.7 [Graphics window], page 28, for more information.

‘Register’

Enter your name in the gadget labeled ’Your Name’. If you have registered, enter your registration key number in the ’Registration Key’ gadget, otherwise leave it blank. See Section 7.8 [Register window], page 29, for more information.

Once you have your configuration set, click on the ’Save’ gadget of the ’Config’ window to save it to disk. If you think you messed up somewhere and want to start again from scratch, click on ’Cancel’. If you want to try out your configuration before you save it, click on ’Use’.

5 Using AmiQWK

5.1 Starting AmiQWK

Before you use ‘AmiQWK’, You will need to download a QWKMail packet from your favorite BBS. Consult the online documentation for the BBS QWK mailer (or, if all else fails, the system operator) for instructions on how to download a QWKMail packet. ‘AmiQWK’ will run without any packets, but you won’t be able to do much with it.

TIP: Do NOT use Zmodem Resume mode when downloading QWKMail packets.

Since most BBSes use the same filename for packets every time, if you already have a QWK packet in your downloads directory your term program will think you are trying to resume an aborted transfer, and the new packet will be appended to the old one instead of creating a new file.

Once you have a QWKMail packet, start ‘AmiQWK’ by double clicking the ‘AmiQWK’ icon or typing ’‘AmiQWK’’ from the CLI. ‘AmiQWK’ will open a custom, public screen (named ’AMIQWK’) with a window titled ’Select Packet’. (If you haven’t registered yet, a window will appear reminding you to register before the ’Select Packet’ window (See Section 5.2 [Select Packet window], page 9.) appears, as well as several other places in the program.) See Section 2.2 [Registration], page 3, for more details about registering ‘AmiQWK’.

5.2 Select Packet window

From the Select Packet window, you can open a QWKMail packet for reading, delete or rename packets, enter the configuration editor, and get information about the program. The Select Packet window has the following gadgets:

‘Registered To’

Name of the registered user. See Section 5.3 [Registered To], page 10.

‘Open Packet. . .’

Open a packet for reading. See Section 5.4 [Open Packet...], page 10.

‘Config AmiQWK. . .’

Open the configuration editor. See Section 7.1 [Main Configuration Window], page 23.

‘About AmiQWK. . .’

Displays version, and copyright information, as well as how to get technical support and where to send your registration fees. See Section 5.6 [About AmiQWK...], page 10.

‘Rename Packet. . .’

Rename or move a packet. See Section 5.7 [Rename Packet...], page 11.

‘HELP!!!. . .’

Get online help. See Section 5.8 [HELP!!!...], page 11.

‘Delete Packet. . .’

Delete a packet. See Section 5.9 [Delete Packet...], page 11.

‘Quit and Exit. . .’

Exit program. See Section 5.10 [Quit and Exit...], page 11.

5.3 Registered To

This read only text gadget displays the name of the person who has registered this copy of ‘AmiQWK’, or ’** UNREGISTERED EVALUATION COPY **’ if ‘AmiQWK’ cannot process the registration key. This gadget is read-only. See Section 2.2 [Registration], page 3, for more information on registering ‘AmiQWK’.

5.4 Open Packet…

When the ’Open Packet…’ gadget is selected, ‘AmiQWK’ opens a file requester for you to select the mail packet to be opened. See Section 5.11 [Opening a packet], page 11, for more information.

5.5 Config AmiQWK…

Clicking on this gadget opens the configuration editor. See Section 7.1 [Main Configuration Window], page 23, for more information.

5.6 About AmiQWK…

This gadget opens a window containing version and copyright information, and where to send registrations, bug reports, suggestions, etc.

5.7 Rename Packet…

This gadget opens a file requester to select a file to be renamed. Once a file is selected, another file requester is opened to select the new file name.

A file can be moved by using the rename function. Just enter a different path in the ’Rename File As…’ file requester. You can even ’Rename’ a file to another volume.

5.8 HELP!!!…

This gadget brings up ‘AmigaGuide’ help for the Packet Selection window. You can get help for most windows by pressing the ’HELP’ key.

5.9 Delete Packet…

This gadget opens a file requester to select a file to be deleted. Once a file is selected another requester will appear to ask for confirmation before deleting the file.

5.10 Quit and Exit…

Click on this gadget to exit the program. Unless ’Ask Before Exit’ in the Operations config window (See Section 7.2 [Operation window], page 24.) is turned off you will be asked for confirmation before ‘AmiQWK’ closes.

5.11 Opening a packet

When ’Open Packet…’ is selected from the ’Select Packet’ window, a file requester will appear to select the packet to open. Once a packet is selected, ‘AmiQWK’ closes the ’Select Packet’ window, extracts the mail packet into the work directory (automatically determining what method was used to compress the packet) as well as the reply packet, if it exists. The extraction procedure takes place in a CLI window that appears on the ‘AmiQWK’ screen. (assuming you haven’t changed the spawn window specifications in the Operations configuration window. (See Section 7.2 [Operation window], page 24.))

Once the packet is extracted, ‘AmiQWK’ will scan the packet, building lists of conferences and messages. This may take a while if you are reading a large packet. If index files are not present in the packet or if the number of messages in the messages.dat file does not match the number of message indexes, ‘AmiQWK’ will attempt to generate the indexes itself. (note that the indexes are not saved back to the packet, as ‘AmiQWK’ can generate the indexes very quickly.) Once the packet has been processed, the ’Select item to read’ window will open. See Section 5.12 [Select item to read window], page 12.

5.12 Select item to read window

This window gives important information about the packet and the BBS that it came from, and has several gadgets for selecting various parts of the mail packet for reading. (note that some of the gadgets may be disabled if the corresponding item is not available in the packet.) There is also a menu which can be used to generate requests to add, drop, or reset conferences (assuming that the QWK packer that generated the packet can process these requests, otherwise the menu items will be disabled.)

The following information about the packet and the BBS it came from is displayed on the left part of the window.

‘Packet:’ The filename of the packet (not including path).

‘BBS:’ The name of the BBS that the packet came from.

‘Location:’

The location of the BBS.

‘Sysop:’ The name of the system operator of the BBS.

‘Phone:’ Primary phone number of the BBS.

‘Packet ID:’

Packet identifier for generating reply packets.

‘Time:’ Time that the packet was created.

‘Date:’ Date that the packet was created.

‘User Name:’

Name of the person who downloaded the packet.

And on the right are the following gadgets. Note that some of them may be disabled if they are not available.

‘'Hello' File. . .’

Displays welcome text using the internal text viewer. (See Section 5.13 [Text Viewer], page 14.) Usually is the login and welcome screen from the BBS, but may contain other information.

‘Bulletins. . .’

Display bulletin file(s) using the internal text viewer. (See Section 5.13 [Text Viewer], page 14.) A file requester will be opened to select which file to read. The bulletin files are named ’BLT-x.y’ where ’x’ is the conference the bulletin came from and ’y’ is the actual number of the bulletin. (i.e. BLT-3.1 indicates that it is the first bulletin in conference 3.) Bulletins are usually only included if they have been updated since the last time you called the BBS. Select ’Cancel’ on the file requester to return to the ’Select item to read’ window.

‘News…’ Display the News file using the internal text viewer. (See Section 5.13 [Text Viewer], page 14.) This file can contain anything the sysop chooses. Usually this file, like bulletins, is only included if it has been updated since the last mail packet was downloaded.

‘Read Mail. . .’

Read messages. (See Section 6.1 [Reading Messages], page 16.)

‘New Files. . .’

Displays a list of files recently posted to the BBS using the internal text viewer. (See Section 5.13 [Text Viewer], page 14.)

‘'Goodbye' Text’

Displays logout text using the internal text viewer. (See Section 5.13 [Text Viewer], page 14.) Usually this is the text that the BBS displays to users logging off of the system, however it may contain other information.

‘Close Packet. . .’

Closes the packet currently being read and returns to the Select Packet window. (See Section 5.2 [Select Packet window], page 9.) If there are any replies, they will be packed into a reply packet for uploading.

‘Exit AmiQWK. . .’

Exit program. Unless ’Ask Before Exit’ in the Operations config window is turned off (See Section 7.2 [Operation window], page 24.) you will be asked for confirmation before ‘AmiQWK’ closes. If there are any replies, they will be packed into a reply packet for uploading.

And finally, there is a menu named ’Conference’. It will contain the following menu items, however these items may be disabled if the QWK packer cannot handle these requests.

‘Add Conference’

Add a conference to your list of active conferences. A window will appear listing all conferences that don’t have any messages in them (if they have messages, you must already be subscribed to that conference.) once a conference is selected, you will be asked for confirmation before the add conference request is generated.

‘Drop Conference’

Remove a conference from your list of active conferences. A window will appear listing all conferences. Once a conference is selected, you will be asked for confirmation before the drop conference request is generated.

‘Reset Pointer’

Reset last read pointer within a conference. A window will appear listing all conferences. Once a conference is selected you will be asked to enter a new last read message number. Once a last read message number is selected, you will be asked for confirmation before the reset request is generated.

See Appendix A [Advanced QWK features], page 30, for more details on Adding, Dropping, and Reseting conferences.

5.13 Text Viewer

‘AmiQWK’ has a built-in text viewer for reading text files included in the QWKMail packet, such as the Hello, Goodbye, News, Bulletins, and recent uploads files. ANSI display sequences are filtered out before the text is displayed. (Note that if the text uses ANSI cursor positioning codes, the text may be jumbled. However most ANSI text uses only color codes.)

The following keys control movement in the text and message viewers:

KEYPAD 8, UP ARROW

Scroll up one line of text.

KEYPAD 2, DOWN ARROW

Scroll down one line of text.

KEYPAD 9 Scroll up one page of text.

KEYPAD 3 SPACE BAR

Scroll down one page of text.

KEYPAD 7 Move to start of text.

KEYPAD 1 Move to end of text.

In addition to the above keys, you can move through the current text by using the scroll bar at the bottom of the screen. (unless the scroll bar is disabled in the Operations configuration window. (See Section 7.2 [Operation window], page 24.)

The following keys (and their corresponding menu items in the ’Commands’ menu) also perform functions:

Q Quit reading text and return to the ’Select item to read’ window. See Section 5.12 [Select item to read window], page 12.

P Print the current text. See Section 6.7 [Printing Text and Messages], page 22.

S Save the current text. See Section 6.6 [Saving Text and Messages], page 21.

A ANSI display of the current message. See Section 5.14 [ANSI Viewer], page 15.

5.14 ANSI Viewer

The ANSI viewer will display the current message or text on a virtual terminal, as if it were being viewed on a terminal screen while connected to a BBS. ANSI display sequences allow for changing the color of the text and background, cursor position, etc.

The ANSI viewer will correctly display most IBM-PC ANSI sequences. ANSI is displayed on it’s own custom 16 color screen, the palette is hard-coded to match the default IBM-PC display.

The ANSI viewer uses the same font as is used for normal message text.

When in the ANSI viewer the following functions are available:

P Pause the display. Display will be suspended until another key is pressed.

Q Quit ANSI display and return to message or text reader.

A Display ANSI text again.

0-9 Change display speed. 0 is fastest (no delay), 9 is slowest (max. delay).

6 Reading and Writing Messages

6.1 Reading Messages

When ’Read Mail…’ is selected from the ’Select item to read’ window, (See Section 5.12 [Select item to read window], page 12.) the ’Select item to read’ window is closed and you will enter the message reader.

When you first enter the message reader, the ’Select Conference’ window (See Section 6.4 [Conference Window], page 20.) will appear for you to select the conference you want to start reading from. It will default to the first conference with messages other than the ’Replies’ conference. When you have selected a conference, ‘AmiQWK’ will start displaying messages.

The message reader is divided into two sections. The top section contains header information and several gadgets, and the bottom section contains the message text.

The following header information is present in the header at the top of the screen:

FROM: Sender of the message.

TO: Person the message is addressed to.

SUBJ: Subject of the message.

DATE: Date the message was entered.

TIME: Time the message was entered.

CONFERENCE:

Name and number of the conference the message is in.

MESSAGE: Message number, Number of this message out of the total available in this conference,

and the total number of messages in this conference. this packet.

REPLY TO:

Number of the message that this message is a reply to. (or blank if it isn’t a reply to another message.)

PRIVATE If displayed, this message is tagged as being Private, usually readable only by the sender, receiver, and system operator.

READ If displayed, this message has already been read.

PERSONAL If displayed, this message is addressed to you personally.

In addition, there are six gadgets in the upper right of the screen.

‘Index’ Opens the Message Index window, to select a message to read. See Section 6.5 [Message Index window], page 21.

‘Conf.’ Opens the Conference window, to select which conference to read. See Section 6.4 [Conference Window], page 20.

‘Reply’ Reply to the current message. See Section 6.2 [Entering and re-editing new messages], page 18.

‘Help’ Call up the ‘AmigaGuide’ help for the message viewer.

‘<<’ Read the previous message.

‘>>’ Read the next message.

At the bottom of the screen is the message text. There is a scroll bar at the bottom (unless the scroll bar is disabled in the Operations configuration window.) which can be used to move through the message using the mouse. The following keys can also be used to move through the message.

KEYPAD 8, UP ARROW

Scroll up one line of message.

KEYPAD 2, DOWN ARROW

Scroll down one line of message.

KEYPAD 9 Scroll up one page of message.

KEYPAD 3, SPACE BAR

Scroll down one page of message.

KEYPAD 7 Move to start of message.

KEYPAD 1 Move to end of message.

KEYPAD 4, LEFT ARROW

Read the previous message in the current conference.

KEYPAD 6, RIGHT ARROW, RETURN

Read the next message in the current conference.

SPACE BAR

Move forward 1 page. If the end of the message is passed the next message is read.

The following keys also perform functions:

A ANSI display of the current message. See Section 5.14 [ANSI Viewer], page 15.

C Open Conference window to select a new conference. See Section 6.4 [Conference Window], page 20.

E Re-edit Reply. (only works in ’Replies’ conference.) See Section 6.2 [Entering and re-editing new messages], page 18.

I Open Message Index window to select which message to read in the current conference. See Section 6.5 [Message Index window], page 21.

K Kill (delete) the current message. (Only works in ’Replies’ conference.)

N Enter a new message. See Section 6.2 [Entering and re-editing new messages], page 18.

P Print the current message. See Section 6.7 [Printing Text and Messages], page 22.

Q, ESC Exit message reader and return to the ’Select item to read’ window. See Section 5.12 [Select item to read window], page 12.

R Reply to the current message. See Section 6.2 [Entering and re-editing new messages], page 18.

S Save the current message to disk. See Section 6.6 [Saving Text and Messages], page 21.

T Add a new tagline to the tagline file. See Section 6.8 [New Tagline window], page 22.

6.2 Entering and re-editing new messages

When you select ’Reply’ or ’New message’ when reading messages, the ’New Message’ window is opened for information about the message. (such as who it is to, subject, etc.) and your editor is opened for you to enter your your message. (Note that you can alter the order the order in which the new message window is opened. You can configure ‘AmiQWK’ to edit the message first then open the new message window, Open the new message window then edit the message, or open the new message window both before and after the message is edited (the default). See Section 7.4 [Replies window], page 26, for more information.)

If you are replying to an existing message, the original message will be quoted in the new message text, with the quote prefix you selected in the Replies configuration window prefixed to each line. See Section 7.4 [Replies window], page 26. When re-editing messages the text will be sent to the editor as-is.

‘To’ This is the person the message is addressed to. If you are replying to an existing message, this will default to the sender of the message you are replying to. If it is a new message, this field will be blank. This field is limited to 25 characters.

‘From’ This is normally your name. You can change this field, however most QWKMail systems will not allow messages to be posted unless this field matches your name. This field is limited to 25 characters due to constraints of the QWKMail format.

‘Subject’ This is the subject of the message. If you are replying to an existing message this field will default to the subject of the original message. (with an optional ’RE:’ prefix, provided that ’RE: prefix for reply subject’ is turned on in the ’Replies’ configuration window and the subject doesn’t already have a ’RE:’ prefix.) Otherwise this field is left blank so you can enter whatever you want here. This field is limited to 25 characters due to constraints of the QWKMail format.

‘Conference’

This is the conference you want the message to be posted to. It defaults to the conference you are currently reading messages in. You can enter the conference number in the integer string gadget, or you can click on the ’Conference’ button and the Conference window will appear for you to select a conference. (See Section 6.4 [Conference Window], page 20.) The name of the conference will appear to the right of the conference number gadget. ‘AmiQWK’ will check the selected conference number against a list of conferences to make sure you have selected a valid conference.

‘Public Status’

This is a cycle gadget that can be either set to ’Public’ or ’Private’. A ’public’ message will be visible to everyone, while a ’Private’ message is usually only visible to the sender, the person who it is addressed to, and the system operator. Note that not all conferences support private messages.

‘Return Receipt’

Tells the BBS to generate a return receipt message when this message is received. This function is available only if it is supported by the QWK mailer. Note that asking for a ’Return Receipt’ will prefix the Subject field with ’RRR’.

‘Tagline’ This is a short one-line message that is appended to the end of your message. You can select a tagline from the list gadget, or type one into the string gadget below it. (If you enter a new tagline, it will be appended to your tagline file for future use.) Taglines are disabled when re-editing messages.

‘Select Another Tagline’

If you don’t like the tagline that ‘AmiQWK’ has selected for you, click on this gadget and a new one will be selected.

‘OK’ Accept the current settings.

‘Cancel’ Cancel the new message.

6.3 Writing messages and BBS Etiquette

When writing a message, Keep in mind that others will have to read it. There are certain things that you can do to

• Avoid typing in all caps. Messages that are in all caps can be difficult to read on some monitors.

• When entering a reply, quote only the relevant information.

• Try to keep your paragraphs short, and insert a blank line between paragraphs.

• Try to keep on the subject and don’t post messages to areas where they don’t belong.

Also remember that many BBSes are connected to BBS networks such as Fidonet and Usenet. Additional rules may apply to network message areas, such as high and low ascii characters (ascii 0-31 and 128-255) and ANSI escape sequences may not be permitted, etc. If you are in doubt as to what is and isn’t permitted in a certain conference, ask the system operator or the network conference moderator for guidelines.

6.4 Conference Window

The Conference window is used to select a conference. It consists of a list gadget listing all available conferences, and an OK and a cancel gadget.

‘Conference List’

Lists all available conferences by name and number, with the total number of messages in each conference. You can select a conference by double-clicking on the desired line of the conference or by selecting a conference and clicking on the ’OK’ gadget.

‘OK’ Accept the currently selected conference and close the Conference window.

‘Cancel’ Closes the Conference window and discards any conference selection.

You can also control the Conference window with the keyboard.

C, Q, ESC

Closes the Conference window and discards any conference selection.

O, ENTER Accept the currently selected conference and close the Conference window.

KEYPAD 8, UP ARROW

Scroll up through the conference list.

KEYPAD 2, DOWN ARROW

Scroll down through the conference list.

6.5 Message Index window

The Message Index window is used to select a message from within a conference. It consists of a list gadget listing all available messages, and an OK and a cancel gadget. The operation of the Message Index window is nearly identical to the Conference window.

‘message List’

Lists all available messages, by the To, From, Subject fields. The following flags are present as well: ‘*’ The message is addressed to you personally. ‘R’ This message has been read. ‘P’ This message is marked as being Private. You can select a message by double-clicking on the desired line of the message list, or by picking the desired conference with either the mouse or keyboard and then clicking on the ’OK’ gadget.

‘OK’ Accept the currently selected message and close the Message Index window. ‘Cancel’ Closes the Message Index window and discards any message selection.

You can also control the message window with the keyboard.

C, Q, ESC

Closes the Message Index window and discards any message selection.

O, ENTER Accept the currently selected message and close the Message Index window.

KEYPAD 8, UP ARROW

Scroll up through the message list.

KEYPAD 2, DOWN ARROW

Scroll down through the message list.

6.6 Saving Text and Messages

When you select ’Save’, A file requester will open for you to select the file name to save the text or message to. If the file you selected already exists the new text or message will be appended to the end of the existing file. If you are saving a message the contents of the header will be saved as well.

The default path and/or directory for the saved text can be set in the ’Paths’ window of the config editor. See Section 7.5 [Paths window], page 27.

6.7 Printing Text and Messages

When you select ’Print’, the current text or message will be sent to the printer. If you are printing a message the contents of the header will be printed as well.

6.8 New Tagline window

The New Tagline window is used to enter a new tagline into the taglines list file. It consists of three gadgets:

‘Tagline’ A string gadget to enter the new tagline into.

‘OK’ Close the New Tagline window and save the new tagline to the tagline list file.

‘Cancel’ Close the New Tagline window and discard the new tagline.

One idea for those of you who like to steal taglines. Use a text snapping program such as PowerSnap to cut taglines out of messages using the mouse and then paste them into the New Tagline window.

7 Configuration

7.1 Main Configuration Window

The configuration editor consists of a main configuration window and several sub-windows. The main Configuration window has the following gadgets:

‘Operation’

Opens the Operation configuration window. The Operation window is used to configure the general operation of ‘AmiQWK’. See Section 7.2 [Operation window], page 24.

‘Replies’ Opens the Replies configuration window. The Relies window is used to configure how replies and new messages are processed. See Section 7.4 [Replies window], page 26.

‘Paths’ Opens the Paths configuration window. The Paths window is used to tell ‘AmiQWK’

where it should look for and store files. See Section 7.5 [Paths window], page 27.

‘Archivers’

Opens the Archivers configuration window. The Archivers window is used to specify which compression programs are used and their command line options. See Section 7.6 [Archivers window], page 27.

‘Graphics’

Opens the Graphics configuration window. The Graphics window is used to specify which screen mode the ‘AmiQWK’ screen should open in, the font that should be used for text displays, the screen palette, etc. See Section 7.7 [Graphics window], page 28.

‘Register’

Opens the Registration configuration window. The Register window is used to enter your registration key when you register ‘AmiQWK’. See Section 7.8 [Register window], page 29.

‘Save’ Accepts the current configuration and saves it to disk and return to the ’Select Packet’ window. See Section 5.2 [Select Packet window], page 9.

‘Use’ Accepts the current configuration and returns you to the ’Select Packet’ window See Section 5.2 [Select Packet window], page 9. but does not save the new configuration to disk. If the new configuration is not saved to disk before exiting the program the new configuration will be lost.

‘Cancel’ Discard all changes to the configuration and return to the ’Select Packet’ window. See Section 5.2 [Select Packet window], page 9.

The configuration is saved in a file named ’Config.AQ’ in the same directory that ‘AmiQWK’ is located in.

7.2 Operation window

The Operation window is used to configure the general operation of ‘AmiQWK’.

‘Scroll bar in text window’

If this checkbox gadget is selected, a scroll bar will appear at the bottom of the text and message windows. This scroll bar can be used to move around the displayed text using the mouse.

‘Tab Size’ This slider gadget specifies the number of spaces between tab stops.

‘Ask before Exit’

If this checkbox is selected, ‘AmiQWK’ will ask for confirmation before exiting. Otherwise it will exit without asking first.

‘After Read’

This cycle gadget is used to tell ‘AmiQWK’ what to do with a packet after you read it. It can be set to:

‘No Action’

Nothing is done to the packet after you read it.

‘Rename Packet’

You will be presented with a file requester to change the name of the packet after you read it.

‘Delete Packet.’

You will be asked if you want to delete the packet after you read it.

‘Ask’ You will be presented with a requester asking if you want to Rename,

Delete, or take no action at all on the packet after you have read it.

‘Packet Search Pattern Match’

This is the pattern that ‘AmiQWK’ should use when looking for mail packets. It should normally be something like ’#?.Q#?’ as packet names normally end with ’.QWK’ and some term programs may change the extension if a duplicate file is downloaded.

‘Editor’ This is the command line used to start your editor for editing messages. Be sure to specify any command line switches you need here as well. The file name will be appended to the end of this string. Keep in mind that your text editor must NOT detach itself from the process that started it. Here are some example editor strings for various editors.

CygnusEd c:ed -sticky (See Section 7.3 [CygnusEd Notes], page 25.)

TurboText
          c:ttx WAIT
Ed        c:ed
MEmacs    sys:tools/memacs
GoldEd    c:ed sticky

If you are using ‘CygnusEd’, See Section 7.3 [CygnusEd Notes], page 25.

‘Workbench to front on edit’

If this checkbox is selected, ‘AmiQWK’ will pull the workbench screen to front when the editor is started. This is useful if you are using an editor that opens on the workbench screen but does not pull it to the front for you, such as the ’ed’ program that comes with Workbench.

‘Spawn Window’

This string gadget is used to set the console specifications for the spawn window for extracting packets and packing replies. See your AmigaDos manual for more information on console specifications. If you want the spawn window to open on the ‘AmiQWK’ screen, make sure that ’SCREENAMIQWK’ appears somewhere on this line.

‘Requesters’

Allows you to choose between ASL and ‘ReqTools’ file, font, and screenmode requesters. (note that ‘AmiQWK’ will use the ‘ReqTools’ screenmode requester if you are using version 37 of ASL.library, because ASL.library v37 does not have a screenmode selector.)

‘Use’ Use the current settings and return to the main configuration window. See Section 7.1 [Main Configuration Window], page 23.

‘Cancel’ Discard changes and return to the main configuration window. See Section 7.1 [Main Configuration Window], page 23.

7.3 CygnusEd Notes

If you are using ‘CygnusEd’ version 2.12 or earlier, you may have trouble getting ‘CygnusEd’ to open without detaching itself from it’s parent process due to a bug in the ‘ED’ activator program. If you are using ‘CygnusEd’ 2.12, there are three options open to you.

• You can upgrade to a newer version of ‘CygnusEd’.

• You can get a third party replacement for the ’‘ED’’ activator program. There are at least two in the public domain. At least one of them will be available from the ‘AmiQWK’ support BBS at all times. See Appendix C [Technical Support], page 32.

• You can use the included ‘CygnusEd’ starter script. (filename ’StartCed’, set the ’Editor’ gadget in the ’Operation’ window of the config editor to ’‘StartCed’’ if you use this script. You may also need to specify the full path to the ‘StartCed’ script.)

For best results, you will probably want to ’hot-start’ ‘CygnusEd’ from the ‘s:user-startup’ file. See your ‘CygnusEd’ program for details on ’hot-starting’.

7.4 Replies window

‘Edit Header’

This cycle gadget has three settings. It controls when the message editor and the ’New Message’ window are opened.

‘Before Edit’

Edit message header in the ’New Message’ window before editing the message text.

‘After Edit’

Edit message header in the ’New Message’ window after editing the message text.

‘Before and After Edit’

Edit message header in the ’New Message’ window both before and after editing the message text.

‘Quote Prefix’

This string is prefixed before each line of quoted text when replying to a message.

‘Header Info in replies’

If this checkbox is selected, the information in the message header will be included in the message text of a reply. Has no effect on new messages. (i.e. not a reply to an existing message.)

‘RE: prefix for reply subject’

If this checkbox is selected, the subject field of the new message will be prefixed with ’RE:’ unless it already has a ’RE:’ prefix.

‘Trunc/Wrap reply text’

Specifies if quoted text in a reply is to be truncated or word wrapped.

‘Col.’ Specifies column that quoted text in a reply is to be truncated or word wrapped at. ‘Use’ Use the current settings and return to the main configuration window. See Section 7.1

[Main Configuration Window], page 23.

‘Cancel’ Discard changes and return to the main configuration window. See Section 7.1 [Main Configuration Window], page 23.

7.5 Paths window

The paths window is used to tell ‘AmiQWK’ where it should look for and store files. To the right of each path string gadget is a ’pickfile’ button, which will open a file requester to select the desired path. You can either type the path into the string gadget or select the path from the file requester.

‘Packet Directory’

Tells ‘AmiQWK’ where it should look for mail packets. This should be set to the downloads directory your terminal program uses.

‘Reply Directory’

Tells ‘AmiQWK’ where it should store reply packets. This should be set to the uploads directory your terminal program uses.

‘Work Directory’

Tells ‘AmiQWK’ where it should create it’s work directory. If you have enough memory, this should be set to ‘RAM:’ for best performance. If this field is left blank, ‘AmiQWK’ will create it’s work directory in the directory in the same directory as the ‘AmiQWK’ executable file.

‘Saved Text’

Default path and/or filename for saved text. You can enter a complete path with filename, just a filename (the files will be saved to the ‘AmiQWK’ directory unless you specify a path when you save messages or text files) or just a path (you will need to make sure that the path name ends with a ’:’ or a ’/’, otherwise ‘AmiQWK’ will think that the last part of the path is actually a file name.)

‘Use’ Use the current settings and return to the main configuration window. See Section 7.1 [Main Configuration Window], page 23.

‘Cancel’ Discard changes and return to the main configuration window. See Section 7.1 [Main Configuration Window], page 23.

7.6 Archivers window

The Archivers window is used to specify the command line strings for the file compression/extraction programs. Currently ‘AmiQWK’ supports ‘LHA/LHArc’, ‘ZIP’, ‘ZOO’, and ‘ARC’ compression. The filename and destination directory will be appended to the command line strings.

‘Use’ Use the current settings and return to the main configuration window. See Section 7.1 [Main Configuration Window], page 23.

‘Cancel’ Discard changes and return to the main configuration window. See Section 7.1 [Main Configuration Window], page 23.

7.7 Graphics window

The Graphics window is used to specify the general appearance of ‘AmiQWK’.

‘Screen Mode’

The name of the screen mode that ‘AmiQWK’ opens on. (Read only, use the ’Select Screen Mode’ gadget to select a screen mode.)

‘Width’ The width of the ‘AmiQWK’ screen in pixels. (Read only)

‘Height’ The height of the ‘AmiQWK’ screen in pixels. (Read only)

‘Select Screen Mode’

Select a screen mode for ‘AmiQWK’ using the ASL or ‘ReqTools’ screenmode requester. (If you are using version 37 of ASL.library, the ‘ReqTools’ screenmode requester will be used.) You can also specify the number of colors the ‘AmiQWK’ screen uses. (I recommend using 4 colors. At this time ‘AmiQWK’ does not use more than four colors except for the ANSI display screen, which is hard coded to use a 16 color screen.)

‘Shanghai’

If this checkmark is selected, the ‘AmiQWK’ screen will be opened in ’Shanghai’ mode. If this option is set, older applications that would normaly open on the Workbench screen instead open on the default public screen, which may or may not be the Workbench screen.

‘PopPubScreen’

If this checkmark is selected, the ‘AmiQWK’ screen will be opened in ’PopPubScreen’ mode. This will cause the AmiQWK screen to ’pop’ to the front whenever a window is opened on the AmiQWK screen.

‘Edit Palette’

Change the color palette ‘AmiQWK’ uses using the ‘ReqTools’ palette requester.

‘Select Font’

Select the font ‘AmiQWK’ uses for message text and the ANSI display, using the ASL or ‘ReqTools’ font requester. The selected font and size will appear to the right of the ’Select Font’ gadget. I recommend using an IBM-PC style font as many packets will contain characters that only appear properly when viewed with an IBM-PC style font.

‘Use’ Use the current settings and return to the main configuration window. See Section 7.1 [Main Configuration Window], page 23.

‘Cancel’ Discard changes and return to the main configuration window. See Section 7.1 [Main Configuration Window], page 23.

7.8 Register window

The Register window is used generate a registration keyfile. When you register you will be given a 16 digit registration number which should be entered into this window.

‘Your Name’

Enter your name into this gadget. It must be entered exactly as it was specified on your registration form as the registration key is matched against this field.

‘Registration Key’

Enter your 16 digit registration number in this gadget.

‘Status’ Registration status. Will either be a registration serial number or ’** UNREGIS-TERED EVALUATION COPY **’ if a valid registration key is not present.

‘Use’ Save your registration key to disk and return to the main configuration window. See Section 7.1 [Main Configuration Window], page 23.

‘Cancel’ Discard changes and return to the main configuration window. See Section 7.1 [Main Configuration Window], page 23.

Appendix A Advanced QWK features

Some BBSes allow for adding, dropping, and resetting of conferences, downloading files attached to messages, and other functions by sending messages to the BBS via a QWKmail packet. If the BBS supports these features a file named DOOR.ID will be included in the QWKMail packet that will tell ‘AmiQWK’ what functions are available and how to process a request.

For example, if you had a packet from a Wildcat! BBS and had ‘AmiQWK’ generate a request to add conference #25, ‘AmiQWK’ would generate a message with the following header information:

FROM: <your name> TO: TOMCAT SUBJ: ADD CONFERENCE 25

If you decide to abort the request, just enter the ’Replies’ conference and ’Kill’ the request message. (See Section 6.1 [Reading Messages], page 16.)

Appendix B The Future of AmiQWK

‘AmiQWK III’will be a complete rewrite of the existing code. Some of the features I plan on adding are:

• A whole new multi-window, multi-packet environment.

− Any number of packets can be opened at a time, limited only by available memory.

• Database of services − New messages can be created even if no packet is present. − Replies can be sent to a different service than the packet is from.

• Sorting − Conferences by Name or Number. − Messages by To, From, Subject, Message Number, etc.

• String search for message by to:, from:, subject:, message text.

• Improvement of quote system.

− Quote prefix may contain initials, first name of the person who’se text is quoted, etc.

• Tagline manager − Edit, Delete, Sort, Find tagline functions. − multiple tagline files can be maintained. (such as a file containing only Star Trek quotes,

etc.) − A true ’Steal Tagline’ function.

• Database of BBS and packet information − Keep a record of which packets have been read, etc. − Keep record of conferences and other information about the packets a BBS generates, so

that a reply packet can be generated even if no packet exists for the BBS. − Keep a address book of people you regularly send messages to, including network addresses,

phone and fax numbers, etc. − Keep a ’Twit’ list. Ignore messages with certain topics or from/to persons listed.

• ’Blue Wave’ and ’QWK-II’ packet support.

• Possible UUCP, TCP/IP, Fidonet, and FAX interfaces.

• UUEncode/UUDecode support.

• PGP (Pretty Good Privacy) support.

• ’Text Variables’ for inserting information into text, such as who the message is to/from, etc.

• ARexx Interface

If there is a feature you would like to see in a future version, let me know about it, even if I listed it already. If I get a lot of requests I may move it to a higher position on the list.

Remember, The future of ‘AmiQWK’ depends on YOU! Send in your registration as soon as possible.

Appendix C Technical Support

Technical support for ‘AmiQWK’ is available through the following:

ELECTRONIC MAIL

INTERNET jim dawson@hguild.com

I also monitor the USENET comp.sys.amiga.datacomm, alt.usenet.offlinereader and comp.sys.amiga.programmer newsgroups.

PORTAL jimdawson (or jimdawson@cup.portal.com via internet.)

FIDONET 1:2290/42.0 Note that at this time I do not monitor any Fidonet echoes that discussions of AmiQWK would be relevant to.

POSTAL MAIL

Jim Dawson

509 S. Madison St. Marion, IL. 62959 USA

BBS The Hackers Guild II BBS (618)997-2591 300-14,400 BPS USR HST/v.32/v.32bis. Full Duplex, 8 data bits, 1 stop bit, no parity. 24 Hrs, 7 Days. (Ex. Fidonet Zone Mail Hour.) You can download from the AmiQWK file area on your first call.

For fastest and most reliable way to get technical support is to call The Hackers Guild II BBS directly. I recommend sending email to my Portal account only if all other methods fail, as I only call Portal once or twice a week.

When asking for help, please give as complete description of your problem as possible, as well as a complete description of your system and any software that might also be running at the same time as ‘AmiQWK’. The more information I have the faster I can suggest a solution to your problem.

I will personally post new versions of AmiQWK to The Hackers Guild II BBS and Portal, and new versions will be submitted to Fred Fish for inclusion in his library of freely distributable software. Users of Fidonet mailers can FREQ the latest version of AmiQWK by requesting ’AMIQWK’ from 1:2290/42.0 (or 618-997-2591 if you don’t have a Fidonet nodelist.) I may also post new versions of AmiQWK to various FTP sites, including the Aminet sites.

Appendix D Acknowledgments

Special thanks go to:

‘Jefferey Foy’

For information on the QWKMail format and MSBIN-IEEE conversion routines.

‘Greg Hewgill’

For specifications of the DOOR.ID file.

‘Patrick Y. Lee’

For his excellent document on the QWKMail format.

‘Marcus Trisdale’

For his ANSI display code.

‘Jan van den Baard’

For his GadToolsBox GUI builder program.

‘Nico François’
          For ‘reqtools.library’
‘Commodore Business Machines’

for ‘AmigaGuide.library’ and for making the Amiga computer.

‘Robert J Chassell and Richard M. Stallman’

for the Texinfo documentation format.

‘Reinhard Spisser and Sebastiano Vigna’
          for ‘MakeGuide’.
‘Chris Hurley, Guy Bartoli, Dave Birdsong, Dennis Jann, Mike Soucy,’
‘Kipling Cooper, Steve Stewart, and Matt Weatherford’

for Beta Testing ‘AmiQWK’

And of course, everyone who has sent in registrations.

Index 35

Index

(Index is empty)

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 About this manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 What is QWKMail? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 What is AmiQWK II?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Legal Stuff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2.1 About Shareware. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Registration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.4 Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.5 Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3 System Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

3.1 Hardware Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 Software Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

4 Installation and Quick Setup . . . . . . . . . . . . . . . . . . . . . . . . 7 5 Using AmiQWK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

5.1 Starting AmiQWK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5.2 Select Packet window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5.3 Registered To . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.4 Open Packet. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.5 Config AmiQWK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.6 About AmiQWK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.7 Rename Packet. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.8 HELP!!!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.9 Delete Packet. . .. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.10 Quit and Exit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.11 Opening a packet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.12 Select item to read window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 5.13 Text Viewer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5.14 ANSI Viewer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

6 Reading and Writing Messages . . . . . . . . . . . . . . . . . . . . 16

6.1 Reading Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 6.2 Entering and re-editing new messages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.3 Writing messages and BBS Etiquette . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.4 Conference Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.5 Message Index window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 6.6 Saving Text and Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 6.7 Printing Text and Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 6.8 New Tagline window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

7 Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

7.1 Main Configuration Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 7.2 Operation window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 7.3 CygnusEd Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 7.4 Replies window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 7.5 Paths window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7.6 Archivers window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7.7 Graphics window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 7.8 Register window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

Appendix A Advanced QWK features. . . . . . . . . . . . . . . 30

Appendix B The Future of AmiQWK . . . . . . . . . . . . . . . 31

Appendix C Technical Support . . . . . . . . . . . . . . . . . . . . . . 33

Appendix D Acknowledgments. . . . . . . . . . . . . . . . . . . . . . . 34

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35